OpenStudy (korosh23):

Exponential Equations Not Requiring Logarithms solve for p (2^-3p) . (2^2p) = (2^2p)

OpenStudy (korosh23):

The way I think to solve this is this way: Since the base is the same. -3p + 2p = 2p -p= 2p How is this possible? What should I write for the answer:?

OpenStudy (unklerhaukus):

yes\[%(2^{-3p}) \cdot (2^{2p}) = (2^{2p})\\ %2^{-3p+2p}=2^{2p}\\ %2^{-p}=2^{2p}\\ -p =2p \] solve for p

OpenStudy (korosh23):

2p +p = 0 3p = 0 p = 0/3 = 0

OpenStudy (unklerhaukus):

yeah, and you can check it by substituting p=0 into eh original equation 1 * 1 = 1

OpenStudy (korosh23):

Right :)

OpenStudy (korosh23):

Thank you for clarifying it.